OpenStudy (anonymous):

I BET NONE OF YOU GUYS CANT ANSWER THIS QUESTION ! Fred has a savings account balance of $1,728.57. The interest rate on the account is 3.4% compounded daily. If he opened the account nine years ago, what was the value of his initial deposit?

OpenStudy (anonymous):

told you ..

OpenStudy (dumbcow):

is this a challange...hmm cant be that hard, banks do it every day

OpenStudy (dumbcow):

any financial calculator will do it for you in seconds...but i wil show the math \[1728.57 = X(1+\frac{.034}{360})^{360*9}\] \[1728.57 = X(1.00009444)^{3240}\] \[X = 1272.91\]